### v2

@@ 4.5.3 Identification of Rule Counterparts
The third criterion evaluates the number of generated rules that have a semantically corresponding rule in the gold standard. Following the method defined in [8], a rule is considered a counterpart if there is a semantic match in its consequent with a rule in the manually curated set of the gold standard. In their evaluation, $33$ out of $36$ gold standard rules had counterparts, resulting in a precision of $80.49\%$ and a recall of $91.67\%$. Our results are presented in Table 3. The best precision, $93.33\%$, is achieved with a fine-tuned version of GPT-4o, whereas the best recall, $84.62\%$, with GPT-o3 in the baseline setting.
Table 3: Precision and Recall for Counterpart Identification
Experiment Model
Generated
Rules
Matched
Rules
Precision Recall F1
Results from [8] $80.49\%$ $\mathbf{91.67\%}$ $0.86$
Simple CoI (Section 4.1) o3 $57$ $44$ $77.19\%$ $84.62\%$ $0.81$
o4-mini $51$ $39$ $76.47\%$ $75.00\%$ $0.76$
DeepSeek-R1 $47$ $42$ $89.36\%$ $80.77\%$ $0.85$
GPT-4o $47$ $42$ $89.36\%$ $80.77\%$ $0.85$
Single Interaction (Section 4.2.3) o4-mini $47$ $41$ $87.23\%$ $78.85\%$ $0.83$
DeepSeek-R1 $31$ $27$ $87.10\%$ $51.92\%$ $0.65$
GPT-4o $36$ $30$ $83.33\%$ $57.69\%$ $0.68$
Fine-tuned (Section 4.3) GPT-4o (Config 1, Epoch 1) $33$ $30$ $90.91\%$ $57.69\%$ $0.71$
GPT-4o (Config 2, Epoch 1) $45$ $41$ $91.11\%$ $78.85\%$ $0.85$
GPT-4o (Config 3, Epoch 1) $45$ $42$ $\mathbf{93.33\%}$ $7\% 0.87\mathbf{0.87}$
Two-Stage (Section 4.4) DeepSeek-R1 $\rightarrow$ DeepSeek-R1 $44$ $36$ $81.82\%$ $69.23\%$ $0.75$
DeepSeek-R1 $\rightarrow$ GPT-4o $44$ $36$ $81.82\%$ $69.23\%$ $0.75$
@@ 4.5.4 Evaluation of Full Rule Correspondence
Finally, we assess the degree of full correspondence, where a rule in the generated set semantically matches the gold standard in both antecedents and consequent. Dragoni et al. report that $24$ rules in their generated set fully matched semantically their counterparts in the gold standard, with a resulting precision of $58.54\%$ and recall of $66.67\%$. The corresponding results from our evaluation are displayed in Table 4. The best precision ($80.56\%$) is achieved with a fine-tuned version of GPT-4o, while the best recall value ($69.23\%$) is reached by multiple approaches simultaneously.
Table 4: Precision and Recall for Full Rule Correspondence
Experiment Model
Generated
Rules
Matched
Rules
Precision Recall F1
Results from [8] $58.54\%$ $66.67\%$ $0.62$
Simple CoI (Section 4.1) o3 $57$ $36$ $63.16\%$ $\mathbf{69.23\%}$ $0.66$
o4-mini $51$ $30$ $58.82\%$ $57.69\%$ $0.58$
DeepSeek-R1 $47$ $31$ $65.96\%$ $59.62\%$ $0.63$
GPT-4o $47$ $32$ $68.09\%$ $61.54\%$ $0.65$
Single Interaction (Section 4.2.3) o4-mini $47$ $36$ $76.60\%$ $\mathbf{69.23\%}$ $0.73$
DeepSeek-R1 $31$ $23$ $74.19\%$ $44.23\%$ $0.55$
GPT-4o $36$ $29$ $\mathbf{80.56\%}$ $55.57\%$ $0.66$
Fine-tuned (Section 4.3) GPT-4o (Config 1, Epoch 1) $33$ $24$ $72.73\%$ $46.15\%$ $0.56$
GPT-4o (Config 2, Epoch 1) $45$ $36$ $80.00\%$ $\mathbf{69.23\%}$ $\mathbf{0.74}$
GPT-4o (Config 3, Epoch 1) $45$ $35$ $77.78\%$ $67.31\%$ $0.72$
Two-Stage (Section 4.4) DeepSeek-R1 $44$ $27$ $61.36\%$ $51.92\%$ $0.56$
GPT-4o $44$ $29$ $65.91\%$ $55.77\%$ $0.60$
@@ 4.5.5 Limitations of the Compared Evaluation Metrics
customerDissatisfiedTimeframe(X) $\Rightarrow$
    [O] informInternalPrioritisation(X)
customerDissatisfiedTimeframe(X) $\Rightarrow$
    [O] informInternalEscalationProcess(X)
escalation(X), internalPrioritisation(X) $\Rightarrow$
    [O] informExternalDisputeResolution(X)
In contrast, some LLMs produced the following additional formalizations:
complaint(X), consumerRequestsUrgent(X) $\Rightarrow$
    [O] informInternalPrioritisation(X)
complaint(X), consumerRequestsUrgent(X) $\Rightarrow$
    [O] informInternalEscalation(X)
@@ 5.1 Legal Interpretation
A major challenge for the formal encoding of legal documents is that each encoding is an interpretation, and the gold standard should correspond to the authentic interpretation. However, in some jurisdictions, it will not be possible to have a true gold standard. The gold standard would correspond to the authentic interpretation of the legal provision, and the only authority able to provide an authentic interpretation is the judiciary. Moreover, this is possible only for cases disputed in court, and it would be limited to the provisions effectively used in the legal proceeding. The second issue is that a legal interpretation depends on the understanding of the legal intent, legal context and the encoding style of the coders. [28] reports on an empirical experiment where three (experienced) coders were asked to model in DDL the same set of legal provisions (from the Australian Copyright Act). The experiment had two phases; in the first phase, the coders did the encoding fully independently. In the second phase, the coders agreed on a common set of terms and then encoded independently the provisions as rules. In the first phase, the degree of agreement varied from 0% to 10% for terms, and 0% of rules using a perfect match, and around 50% for terms and 3% on rules with a semantics correspondence. In phase two, the term agreement was between 30% and 56% for the full correspondence and 85% for semantic correspondence; the rule similarity ranged from 10% to 30% for full correspondence and 26% to 53% with semantic correspondence.
@@ 5.2 Inter-Paragraph References
A common strategy for managing the complexity of legal documents is the use of references, which may be internal (linking to other sections within the same document) or external (referring to other documents) [29]. Accordingly, the dataset used in this study included occasional references between different legal paragraphs. For example, law snippet 8.2.1.a.xiv mandates that a complaint must only be closed with the consent of the Consumer or former Customer or if clauses 8.2.1(c),(d) or (e) below have been complied with.
@@
This problem persisted even when using the methodology presented in Section 4.2. This limitation suggests that prompt engineering alone is insufficient to fully address the challenge of reference resolution. Instead, it requires the incorporation of additional procedural components into the methodology. One possible solution is a refinement phase following the initial generation process, designed to ensure semantic coherence across references.
5.3 Atom Reuse Across Legal Snippets
Effective formalization of legal texts requires the consistent reuse of atoms across rules; otherwise, reasoning with the resulting rule set containing many redundant atoms would not be useful. However, the experimental results revealed that the LLMs rarely reused atoms across different law snippets in the current setup, except for complaint(X) which was explicitly required in the prompt. For instance, while the gold standard consistently employed the atom resolvable15Days(X), most LLMs generated varying alternatives such as resolveIn15Days(X), resolvedBy15Days(X) and cannotResolveIn15Days(X) across different law snippets. Although OpenAI’s newer reasoning models (e.g., o3 and o4-mini) have shown success in various evaluations, this issue is particularly significant in their outputs, as they produced up to 96 atoms compared to the 69 present in the gold standard.
Possible strategies to address this problem include:
(i) As noted in Section 4.2.2, simply providing LLMs with all previously generated atoms did not improve reuse; instead, it increased hallucinations. A more effective approach may involve identifying the most relevant atoms in advance and selectively providing only those. Alternatively, previously generated atoms may be supplied exclusively to the initial atom extraction phase within a multi-phase pipeline, while subsequent phases focus on generating logically coherent DDL rules, potentially correcting earlier hallucinations.
(ii) We could introduce an intermediate step between atom generation and DDL rule formalization, in which similar atoms could be clustered and evaluated by an LLM to determine whether they should be merged, thereby promoting consistency and reuse.
@@ 6 Conclusion and Future Work
This paper showed that LLMs can be leveraged to transform legal norms into formal DDL rules with substantial fidelity, and with performance similar to the approach proposed in [8]. Our evaluations confirm that prompt engineering – especially few-shot learning with Chain-of-Instructions – can significantly improve the semantic precision of extracted rules. Fine-tuning offers benefits but risks overfitting, while multi-stage pipelines are promising but sensitive to the quality of initial atom extraction.
@@
Acknowledgements.
This work has been partially funded by the Vienna Science and Technology Fund (WWTF) [Grant ID: 10.47379/ICT23030].

### v3

@@ 4.5.3 Identification of Rule Counterparts
The third criterion evaluates the number of generated rules that have a semantically corresponding rule in the gold standard. Following the method defined in [DBLP:conf/aicol/DragoniVRG17], a rule is considered a counterpart if there is a semantic match in its consequent with a rule in the manually curated set of the gold standard. In their evaluation, $33$ out of $41$ generated rules had counterparts among the reported $36$ gold standard rules, resulting in a precision of $80.49\%$ and a recall of $91.67\%$. The metrics evaluation computed with the reported $36$ gold standard rules, the evaluation corrected with the $52$ rules we counted, and the results from our models are all presented in Table 3.
The highest precision (87.18%) was achieved by DeepSeek-R1 (0528) when formalizing all law snippets in a single interaction. The highest recall (82.69%) was attained by GPT-4.1 and o3 in Section 4.1, as well as by using GPT-4.1 as first model in Section 4.4.2. The highest F1-score was obtained with a fine-tuned version of GPT-4.1, as well as by using o3 as model for the initial formalization in Section 4.4.2.
Table 3: Precision, Recall and F1-score for Counterpart Identification
Experiment Model Matched/ Generated Rules Precision Recall F1
Results from [DBLP:conf/aicol/DragoniVRG17] reported 36 rules 33/41 $80.49\%$ $91.67\%$ $0.86$
counted 52 rules1 $80.49\%$ $63.46\%$1 0.710.711
Simple CoI (Section 4.1) GPT-4.1 $43/56$ $76.79\%$ $\mathbf{82.69\%}$ $0.80$
Cl. Sonnet 4 (ET) $41/50$ $82.00\%$ $78.85\%$ $0.80$
DeepSeek-R1 (0528) $36/43$ $83.72\%$ $69.23\%$ $0.76$
o3 $43/57$ $75.44\%$ $\mathbf{82.69\%}$ $0.79$
o4-mini $37/51$ $72.55\%$ $71.15\%$ $0.72$
GPT-4o $39/47$ $82.98\%$ $75.00\%$ $0.79$
Single Interaction (Section 4.2.3) GPT-4.1 $30/37$ $81.08\%$ $57.69\%$ $0.67$
DeepSeek-R1 (0528) $34/39$ $\mathbf{87.18\%}$ $65.38\%$ $0.75$
o4-mini $40/47$ $85.11\%$ $76.92\%$ $0.81$
GPT-4o $29/36$ $80.56\%$ $55.77\%$ $0.66$
Fine-tuned after Epoch 1 (Section 4.3) GPT-4.1 (Config 2) $38/45$ $84.44\%$ $73.08\%$ $0.78$
GPT-4.1 (Config 3) $40/46$ $86.96\%$ $76.92\%$ $\mathbf{0.82}$
GPT-4o (Config 2) $38/45$ $84.44\%$ $73.08\%$ $0.78$
GPT-4o (Config 3) $38/45$ $84.44\%$ $73.08\%$ $0.78$
Refinement Step w. Claude Sonnet 4 (Section 4.4.2) GPT-4.1 $43/56$ $76.79\%$ $\mathbf{82.69\%}$ $0.80$
Cl. Sonnet 4 (ET) $40/50$ $80.00\%$ $76.92\%$ $0.78$
DeepSeek-R1 (0528) $36/43$ $83.72\%$ $69.23\%$ $0.76$
o3 $42/51$ $82.35\%$ $80.77\%$ $\mathbf{0.82}$
  • 1
    Note. We counted 52 rules in the gold standard as opposed to the 36 reported in [DBLP:conf/aicol/DragoniVRG17]. Therefore, we present both sets of metric values: those originally reported and those recalculated based on our count. For the purpose of comparing with our LLM-based approach, we use the recalculated values as reference.
@@ 4.5.4 Evaluation of Full Rule Correspondence
Finally, we assess the degree of full correspondence, where a rule in the generated set semantically matches the gold standard in both antecedents and consequent. Dragoni et al. report that $24$ rules in their generated set fully matched semantically their counterparts in the gold standard, with a resulting precision of $58.54\%$ and recall of $66.67\%$. The corresponding results from our evaluation are displayed in Table 4.
The highest precision ($76.92\%$) was achieved by DeepSeek-R1 (0528) in Section 4.2.3, while the highest recall ($75\%$) was achieved by using GPT-4.1 as first model in Section 4.4.2. Moreover, the highest F1-score ($0.74$) was achieved by refining the output of o3 in a two-stage pipeline.
Table 4: Precision, Recall and F1-score for Full Rule Correspondence
Experiment Model Matched/ Generated Rules Precision Recall F1
Results from [DBLP:conf/aicol/DragoniVRG17] reported 36 rules $24/41$ $58.54\%$ $66.67\%$ $0.62$
counted 52 rules1 $58.54\%$ $46.15\%$1 0.520.521
Simple CoI (Section 4.1) GPT-4.1 $37/56$ $66.07\%$ $71.15\%$ $0.69$
Cl. Sonnet 4 (ET) $36/50$ $72.00\%$ $69.23\%$ $0.71$
DeepSeek-R1 (0528) $31/43$ $72.09\%$ $59.62\%$ $0.65$
o3 $37/57$ $64.91\%$ $71.15\%$ $0.68$
o4-mini $30/51$ $58.82\%$ $57.69\%$ $0.58$
GPT-4o $30/47$ $63.83\%$ $57.69\%$ $0.61$
Single Interaction (Section 4.2.3) GPT-4.1 $28/37$ $75.68\%$ $53.85\%$ $0.63$
DeepSeek-R1 (0528) $30/39$ $\mathbf{76.92\%}$ $57.69\%$ $0.66$
o4-mini $36/47$ $76.60\%$ $69.23\%$ $0.73$
GPT-4o $27/36$ $75.00\%$ $51.92\%$ $0.61$
Fine-tuned after Epoch 1 (Section 4.3) GPT-4.1 (Config 2) $33/45$ $73.33\%$ $63.46\%$ $0.68$
GPT-4.1 (Config 3) $33/46$ $71.74\%$ $63.46\%$ $0.67$
GPT-4o (Config 2) $34/45$ $75.56\%$ $65.38\%$ $0.70$
GPT-4o (Config 3) $33/45$ $73.33\%$ $63.46\%$ $0.68$
Refinement Step w. Claude Sonnet 4 (Section 4.4.2) GPT-4.1 $39/56$ $69.64\%$ $\mathbf{75.00\%}$ $0.72$
Cl. Sonnet 4 (ET) $36/50$ $72.00\%$ $69.23\%$ $0.71$
DeepSeek-R1 (0528) $32/43$ $74.42\%$ $61.54\%$ $0.67$
o3 $38/51$ $74.51\%$ $73.08\%$ $\mathbf{0.74}$
  • 1
    Note that we counted 52 rules in the gold standard as opposed to the 36 reported in [DBLP:conf/aicol/DragoniVRG17]. Therefore, we provide calculations based on both figures.
@@ 4.5.5 Limitations of the Compared Evaluation Metrics
customerDissatisfiedTimeframe(X) $\Rightarrow$ [O] informInternalPrioritisation(X)
customerDissatisfiedTimeframe(X) $\Rightarrow$ [O] informInternalEscalationProcess(X)
escalation(X), internalPrioritisation(X) $\Rightarrow$ [O] informExternalDisputeResolution(X)
However, some LLMs produced the following additional formalizations:
complaint(X), consumerRequestsUrgent(X) $\Rightarrow$ [O] informInternalPrioritisation(X)
complaint(X), consumerRequestsUrgent(X) $\Rightarrow$ [O] informInternalEscalation(X)
@@ 5.1 Legal Interpretation
A major challenge for the formal encoding of legal documents is that each encoding is an interpretation, and the gold standard should correspond to the authentic interpretation. However, in some jurisdictions, it will not be possible to have a true gold standard. The gold standard would correspond to the authentic interpretation of the legal provision, and the only authority able to provide an authentic interpretation is the judiciary. Moreover, this is possible only for cases disputed in court, and it would be limited to the provisions effectively used in the legal proceeding. The second issue is that a legal interpretation depends on the understanding of the legal intent, legal context and the encoding style of the coders. [ail2022:Alice] reports on an empirical experiment where three (experienced) coders were asked to model in DDL the same set of legal provisions (from the Australian Copyright Act). The experiment had two phases; in the first phase, the coders did the encoding fully independently. In the second phase, the coders agreed on a common set of terms and then encoded independently the provisions as rules. In the first phase, the degree of agreement varied from 0% to 10% for terms, and 0% of rules using a perfect match, and around 50% for terms and 3% on rules with a semantics correspondence. In phase two, the term agreement was between 30% and 56% for the full correspondence and 85% for semantic correspondence; the rule similarity ranged from 10% to 30% for full correspondence and 26% to 53% with semantic correspondence.
@@ 5.2 Inter-Paragraph References
A common strategy for managing the complexity of legal documents is the use of references, which may be internal (linking to other sections within the same document) or external (referring to other documents) [DBLP:conf/icail/GovernatoriO21]. Accordingly, the dataset used in this study included occasional references between different legal paragraphs. For example, law snippet 8.2.1.a.xiv mandates that a complaint must only be closed with the consent of the Consumer or former Customer or if clauses 8.2.1(c),(d) or (e) below have been complied with.
@@
This problem persisted even when formalizing all law snippets in a single interaction or when applying a refinement step. This limitation suggests that prompt engineering alone is insufficient to fully address the challenge of reference resolution. Instead, it requires the incorporation of additional procedural components into the methodology. One possible solution is a refinement phase following the initial generation process, designed to ensure semantic coherence across references.
@@ 6 Conclusion and Future Work
This paper demonstrated that LLMs can be effectively leveraged to transform legal norms into formal DDL rules, achieving performance that is in most respects superior to the approach proposed in [DBLP:conf/aicol/DragoniVRG17]. Based on recalculated metrics that reflect the actual number of rules and atoms in the gold standard, our method outperforms Dragoni et al. across nearly all evaluation settings. The only exception concerns precision in the term identification task, where Dragoni et al. attains a higher score.
Our evaluation indicates that prompt engineering – particularly the use of few-shot learning combined with Chain-of-Instructions – yields promising results. By contrast, neither the direct formalization of all law snippets in a single step nor fine-tuning led to measurable performance gains. Moreover, decomposing the task into a two-step pipeline consisting of atom extraction followed by rule generation resulted in a decline in output quality. A modest improvement was, however, observed when employing a two-step pipeline with an additional refinement stage after rule generation. This refinement step contributed, among other aspects, to greater consistency in the use of atom names across law snippets.
@@
[…]
Acknowledgments This work has been partially funded by the Vienna Science and Technology Fund (WWTF) [Grant ID: 10.47379/ICT23030].